OpenStudy (anonymous):

A wooden stage rests on a triangular stilt. The base of the triangle is the same as the height. The area of the stilt is 24.5 square meters. What is the height of the triangle?

OpenStudy (imstuck):

Need help setting it up? Here's a picture! I'm quite the artist....

OpenStudy (imstuck):

The formula for the area of a triangle is A = 1/2 b * h. If the base and the height are equal to one another and they are both unknown, let's call them both x. Filling that into our formula along with the given area you have\[24.5=\frac{ 1 }{ 2 }x \times x\]or\[24.5=\frac{ x ^{2} }{ 2 }\]Can you solve that for x?

OpenStudy (anonymous):

I got 7, is that right?

OpenStudy (anonymous):

That doesn't seem right

OpenStudy (imstuck):

YES!!!! Very good! You go!

OpenStudy (anonymous):

It is?

OpenStudy (anonymous):

Thank you so much

OpenStudy (imstuck):

Yes it is!

OpenStudy (imstuck):

And you're welcome!

OpenStudy (anonymous):

:)